OpenStudy (anonymous):

Sydney needs to earn $42 so she can buy her father a birthday present. She can make $7 per hour working at a local gift shop and $8 per hour babysitting for her neighbors. There are 15 hours available before Sydney's father's birthday where she can either work at the gift shop or babysit. Which system of equations can be used to find how many hours Sydney should spend either working at the gift shop or babysitting so that she will earn enough money to buy her father a birthday present? Let x represent the number of hours spent working at the gift shop and y represent the number of hours spe

OpenStudy (anonymous):

A.$8x + $7y - $42 = 0 x + y - 15 = 0 B.$7x + $8y + $42 = 0 x + y - 15 = 0 C.$7x - $8y - $42 = 0 x + y + 15 = 0 D.$7x + $8y - $42 = 0 x + y - 15 = 0